# Density-Based Long-Range Electrostatic Descriptors for Machine Learning Force Fields

### v1

@@ Abstract
This study presents a long-range descriptor for machine learning force fields (MLFFs) that maintains translational and rotational symmetry, similar to short-range descriptors while being able to incorporate long-range electrostatic interactions. The proposed descriptor is based on an atomic density representation and is structurally similar to classical short-range atom-centered descriptors, making it straightforward to integrate into machine learning schemes. The effectiveness of our model is demonstrated through comparative analysis with the long-distance equivariant (LODE) (Grisafi and Ceriotti, 2019) descriptor. In a toy model with purely electrostatic interactions, our model achieves errors below 0.1%.
The application of our descriptors, in combination with local descriptors representing the atomic density, to materials where monopole-monopole interactions are important such as sodium chloride successfully captures long-range interactions, improving predictive accuracy. The study highlights the limitations of the combined LODE method in materials where intermediate-range effects play a significant role. Our work presents a promising approach to addressing the challenge of incorporating long-range interactions into MLFFs, which enhances predictive accuracy for charged materials to the level of state-of-the-art Message Passing Neural Networks.
@@ I Introduction
To increase the efficiency of MLFFs, many methods rely on a local descriptor that solely represents the atomic environment in the vicinity of an atom within a specified cutoff sphere. These methods assume that all interactions between atoms further apart than the cutoff radius are negligible. While this approach accurately describes many materials and their properties, it neglects long-range interactions such as electrostatics. This is because, in practice, it is impossible to arbitrarily increase the cutoff radius. However, these long-range interactions can be important Kjellander (2018); Guo et al. (2018); Jorn et al. (2013); French et al. (2010), and new methods that include long-range interactions are necessary to improve the predictive power of MLFFs for systems with long-range electrostatics Ko et al. (2021); Unke et al. (2021b); Ambrosetti et al. (2016); Stöhr and Tkatchenko (2019); Unke et al. (2022).
This issue is widely recognized, and various methods have been developed to address it. Recent models have concentrated on specific types of long-range interactions, such as electrostatics, and have introduced correction terms to account for them Grisafi and Ceriotti (2019); Grisafi, Nigam, and Ceriotti (2021); Ko et al. (2021); Unke et al. (2021b); Zhang et al. (2022); Bereau et al. (2018); Yao et al. (2018); Unke and Meuwly (2019); Niblett, Galib, and Limmer (2021); Gao and Remsing (2022). A common approach is to split the energy into distinct parts and model them independently. Then, all contributions are added together to calculate the total energy. Short-range contributions are typically calculated with a local model, while long-range components are modeled with diverse approaches. One can either treat the various energy terms separately and train two machine learning (ML) models, adding up the final energies, or utilize combined descriptors and train a single model.
@@
Another approach to include long-range interactions is to use short-range descriptors, but effectively propagate interactions beyond the distinct cutoff by employing Message Passing Neural Networks (MPNNs) Gilmer et al. (2017); Schütt et al. (2018); Bronstein et al. (2021); Unke et al. (2021b); Batzner et al. (2022); Batatia et al. (2022); Haghighatlari et al. (2022); Gasteiger, Groß, and Günnemann (2022); Takamoto, Izumi, and Li (2022); Chen and Ong (2022); Deng et al. (2023); Merchant et al. (2023); Batatia et al. (2023). Equivariant MPNNs such as NewtonNet Haghighatlari et al. (2022), DimeNet Gasteiger, Groß, and Günnemann (2022), TeaNet Takamoto, Izumi, and Li (2022), NequIP Batzner et al. (2022) and MACE Batatia et al. (2022) offer unique advantages due to their specialized architectures.
NewtonNet is designed to respect Newtonian mechanics, ensuring that the learned forces obey physical laws. It also incorporates directional information from forces, enhancing the model’s accuracy Haghighatlari et al. (2022). DimeNet employs directional message passing, which allows it to capture complex interactions effectively Gasteiger, Groß, and Günnemann (2022). TeaNet uses tensor-based message passing, facilitating the learning of intricate relationships between atomic structures Takamoto, Izumi, and Li (2022).
NequIPBatzner et al. (2022) employs equivariant convolutions of tensorial quantities, resulting in a particularly flexible network topology with many millions of parameters. MACE Batatia et al. (2022) is in most respects a simplified version of NequIP usually relying on only two message-passing layers and largely linear activation functions. This should improve execution speed and learning efficiency.
This work utilizes kernel methods. Our objective is to find a physically meaningful and flexible approach for describing long-range interactions without resorting to a global and non-atom-centered description. We present a new descriptor encoding the atomic density. The atomic density is described similarly to short-range models Jinnouchi, Karsai, and Kresse (2019); Jinnouchi et al. (2019, 2020a, 2020b). However, we implicitly account for all periodic images of all atoms in the supercell by treating the atomic density in reciprocal space.
This approach offers the advantage of a flexible, physics-based descriptor that is atom-centered but also long-ranged. It can be easily combined with a local descriptor, as both have the same mathematical form, albeit in its present implementation it is not yet very performant.
We apply the new approach to a gas of point charges. The descriptor is then compared to the long-distance equivariant (LODE) framework Grisafi and Ceriotti (2019) and the MPNN MACE Batatia et al. (2022) for liquid sodium chloride and zirconia.
@@ II.1 Short-Range Descriptors
from Jinnouchi, Karsai, and Kresse (2019) that are given there in Eq. (18) and (19) are used for building the short-range descriptors. Here, the vector $\textbf{r}_{jk}$ is the vector pointing from atom $j$ to $k$, $\textbf{r}_{k}-\textbf{r}_{j}$. The set of radial $\chi_{nl}$ and angular $Y_{lm}$ basis functions are used to expand the atomic density in order to obtain rotational and translational invariant expansion coefficients Bartók et al. (2010); Bartók, Kondor, and Csányi (2013). The indices $l$, $m$, and $n$ are the angular, momentum, and radial index, respectively. The parameter $\sigma$ is used to broaden the density distribution. $\iota_{l}$ are the modified spherical Bessel functions of the first kind, and $N_{J}$ is the number of atoms of type $J$. The cutoff function $f_{\text{cut}}$ is used to ensure a smooth decay of $h_{nl}$ to zero at the cutoff radius.
@@ II.2 Long-Range Descriptors
The objective is to calculate a similar set of descriptors where the expansion coefficients however represent the atomic density over long ranges and multiple supercells. The same ansatz as in real space is used, where the expansion coefficients are designed to represent the atomic density. As before, the descriptors are obtained by projecting the density onto a product basis of radial and angular functions. However, to avoid introducing a cutoff, the corresponding products are evaluated in reciprocal instead of real space:
@@
where G are momentum vectors consistent with the periodic boundary conditions, and $f_{nlm}(\textbf{G})$ are the basis functions in reciprocal space and the other terms represent the broadened atomic density distribution of atom $k$ of type $J$ with positions $\textbf{r}_{k}$ around atom $j$ with positions $\textbf{r}_{j}$. The central atom $j$ may be of any atom type present in the material. This allows for the inclusion of interactions between atoms of different types in the descriptor.
@@
The long-range expansion coefficients given in Eq. (9) have the same form as the short-range expansion coefficients given in Eq. (1), but with the appropriate choice of basis functions (see next section), they can encode the long-range density. This is due to the calculation in reciprocal space, where not only the position of an atom is considered but also all periodic images of this atom. It is important to note that these long-range descriptors do not require a cutoff, but of course, it is possible to impose a cutoff by using the FT of finite-ranged real-space functions in Eq. (11).
The individual long-range expansion coefficients are then combined into a vector ("long-range descriptor"). We use only two-body expansion coefficients as shown in Eq. (4) and the corresponding vector $\textbf{X}_{j}^{(2)}$. The three-body long-range coefficients are excluded since the objective here is to describe long-range pairwise interactions. Furthermore, we found no improvements when including long-range three-body terms in any of the systems considered here. This might have many reasons, but we believe that three-body interactions are generally shorter-ranged. Consider, for instance, the case of Van der Waals (vdW) interactions: two-body vdW interactions fall off like $1/r^{6}$, and the three-body terms (Axilrod–Teller) fall off like $1/r^{9}$. Furthermore, three-body electrostatic interactions do not exist, as the Coulomb interaction is strictly two-body in nature.
@@ II.3 Long-Range Radial Basis Functions
Spherical harmonics are also used for the angular part of the long-range descriptor. However, spherical Bessel functions cannot be employed to model the infinitely ranged interactions in the radial part. To maintain a fixed spatial resolution, the number of Bessel functions would need to increase with $R_{\text{cut}}$ and therefore go to infinity.
Given the objective of representing electrostatic interactions that decay slowly with $1/r$, where $r$ is the distance between two atoms, it is evident that exponentially decaying functions are a natural choice. This is because
@@
holds. The first line of Eq. (12) is the Laplace transform of $1/r$, while the second line approximates this integral using a quadrature rule. This demonstrates that the integral over all exponentially decaying functions can represent $1/r$, a fact that has been amply used in quantum chemistry to deal with integrals in many-body perturbation theory Häser and Almlöf (1992); Izmaylov and Scuseria (2008); Doser, Lambrecht, and Ochsenfeld (2008); Ayala and Scuseria (1999); Kaltak, Klimes, and Kresse (2014). However, it is necessary to determine the exponents $\zeta_{n}$ and $N_{\text{max}}$. We define $\zeta_{n}$ on a logarithmic mesh where
@@
and $\zeta_{\text{max}}$ and the scaling constant $s$ are hyperparameters determined numerically. We note that for many-body applications a roughly exponential scaling was also found to be optimal, although we had rigorous procedures for determining an optimal scaling Kaltak, Klimes, and Kresse (2014). Those prescriptions are not easily adaptable to machine learning, though.
@@ II.4 The LODE Implementation
Because of the relation shown in Eq. (14), it is relatively easy, with the density-based model presented above, to implement the LODE descriptors as well. The only required step is to multiply the density-based expansion coefficients by the negative Coulomb kernel before summing over the reciprocal grid points G.
As in the original work of Grisafi and Ceriotti for real materials, the LODE descriptors are combined with short-range descriptors as well, using the approach shown in the following section.
@@ II.5 Combination of Short- and Long-Range Descriptors
To apply the long-range models to realistic materials in which both short-range and long-range interactions are present, the final descriptor vectors X must contain information about both types of interaction. By simply appending the short-range descriptors $\textbf{X}_{\text{sr}}$ and long-range descriptors $\textbf{X}_{lr}$ the final descriptor
@@
It should be noted that no additional weighting for the short- and long-range descriptors is introduced here when combining them, as this was found the change the learning efficiency very little.
@@ II.6 The Fitting Process
The fitting process of our model employs the methods of polipy4vasp, the code used in Ref. Schmiedmayer and Kresse (2024). These methods are implemented analogously to those of the Vienna Ab initio Simulation Package (VASP) Kresse and Furthmüller (1996a, b); Kresse and Joubert (1999) which are presented in Jinnouchi, Karsai, and Kresse (2019); Jinnouchi et al. (2019, 2020a).
@@
of local energy contributions $U_{j}$ that depend on the "local" environment around each atom $j$. Kernel regression is used to describe the non-linear dependence of the energy on the descriptors. Either a Gaussian (radial basis function) kernel
@@
In Eq. (18) $\hat{\textbf{X}}$ indicates that the descriptor X is normalized. The index $b$ in both Eq. (17) and Eq. (18) refers to the kernel basis function to which the descriptor of the current central atom $\textbf{X}_{j}$ is compared.
@@
During training the linear system of equations is solved using singular value decomposition, and singular values smaller than a threshold of $10^{-9}$ times the largest singular value are disposed of.
@@ III Results
For LODE the model with only pairwise descriptors and a body order term of $\nu=1$ is used. This facilitates a comparison with the purely radial description of the density-based long-range descriptors. Furthermore, it was found that the use of three-body terms in our case leads to higher errors compared to the use of only two-body terms. This is likely related to the already discussed two-body nature of long-range Coulomb interactions.

### v3

@@ Abstract
This study presents a long-range descriptor for machine learning force fields (MLFFs) that maintains translational and rotational symmetry, similar to short-range descriptors while being able to incorporate long-range electrostatic interactions. The proposed descriptor is based on an atomic density representation and is structurally similar to classical short-range atom-centered descriptors, making it straightforward to integrate into machine learning schemes. The effectiveness of our model is demonstrated through comparative analysis with the long-distance equivariant (LODE) (Grisafi and Ceriotti, 2019) descriptor. In a toy model with purely electrostatic interactions, our model achieves errors below 0.1%, worse than LODE but still very good.
For real materials, we perform tests for liquid NaCl, rock salt NaCl, and solid zirconia. For NaCl, the present descriptors improve on short-range density descriptors, reducing errors by a factor of two to three and coming close to message-passing networks. However, for solid zirconia, no improvements are observed with the present approach, while message-passing networks reduce the error by almost a factor of two to three. Possible shortcomings of the present model are briefly discussed.
@@ I Introduction
To increase the efficiency of MLFFs, many methods rely on local descriptors that solely represent the atomic environment in the vicinity of an atom within a specified cutoff sphere. These methods assume that all interactions between atoms further apart than the cutoff radius are negligible. While this approach accurately describes many materials and their properties, it neglects long-range interactions such as electrostatics. This is because, in practice, it is impossible to increase the cutoff radius arbitrarily. However, these long-range interactions can be important Kjellander (2018); Guo et al. (2018); Jorn et al. (2013); French et al. (2010), and new methods that include long-range interactions are necessary to improve the predictive power of MLFFs for systems with long-range electrostatics Ko et al. (2021); Unke et al. (2021b); Ambrosetti et al. (2016); Stöhr and Tkatchenko (2019); Unke et al. (2022).
This issue is widely recognized, and various methods have been developed to address it. Recent models have concentrated on specific types of long-range interactions, such as electrostatics, and have introduced correction terms to account for them Grisafi and Ceriotti (2019); Grisafi, Nigam, and Ceriotti (2021); Ko et al. (2021); Unke et al. (2021b); Zhang et al. (2022); Bereau et al. (2018); Yao et al. (2018); Unke and Meuwly (2019); Niblett, Galib, and Limmer (2021); Gao and Remsing (2022). A common approach is to split the energy into distinct parts and model them independently. Then, all contributions are added together to calculate the total energy. Short-range contributions are typically calculated using a local model, while long-range components are modeled with diverse approaches. One can either treat the various energy terms separately and train two machine learning (ML) models, adding up the final energies, or utilize combined descriptors and train a single model.
@@
Another approach to include long-range interactions is to use short-range descriptors, but effectively propagate interactions beyond the distinct cutoff by employing Message Passing Neural Networks (MPNNs) Gilmer et al. (2017); Schütt et al. (2018); Bronstein et al. (2021); Unke et al. (2021b); Batzner et al. (2022); Batatia et al. (2022); Haghighatlari et al. (2022); Gasteiger, Groß, and Günnemann (2022); Takamoto, Izumi, and Li (2022); Chen and Ong (2022); Deng et al. (2023); Merchant et al. (2023); Batatia et al. (2023). Other relevant developments are equivariant MPNNs such as NewtonNet Haghighatlari et al. (2022), DimeNet Gasteiger, Groß, and Günnemann (2022), TeaNet Takamoto, Izumi, and Li (2022), NequIP Batzner et al. (2022) and MACE Batatia et al. (2022) that cover not only invariant but also tensorial quantities.
NewtonNet is designed to respect Newtonian mechanics, ensuring the learned forces obey physical laws. It also incorporates directional information from forces, enhancing the model’s accuracy Haghighatlari et al. (2022). DimeNet employs directional message passing, which allows it to capture complex interactions effectively Gasteiger, Groß, and Günnemann (2022). TeaNet uses tensor-based message passing, facilitating the learning of intricate relationships between atomic structures Takamoto, Izumi, and Li (2022).
NequIPBatzner et al. (2022) employs equivariant convolutions of tensorial quantities and multi-level message passing of equivariant features, resulting in a particularly flexible network topology with many millions of parameters. MACE Batatia et al. (2022) is in most respects a simplified version of NequIP usually relying on only two message-passing layers and largely linear activation functions. This should improve execution speed and learning efficiency.
The present work utilizes kernel methods. Our objective is to find a physically meaningful and flexible approach for describing long-range interactions without resorting to a global and non-atom-centered description. Similarly to short-range models Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Jinnouchi, Karsai, and Kresse (2019); Jinnouchi et al. (2019, 2020a, 2020b), the atomic density around each atom is expanded into a suitable set of basis functions. However, we implicitly account for all periodic images of all atoms in the supercell by performing the calculations in reciprocal space and using suitably slowly decaying basis functions.
This approach offers the advantage of a flexible, physics-based descriptor that is atom-centered but also long-ranged. It can be easily combined with a local descriptor, as both have the same mathematical form, albeit in our present implementation the approach is not yet very performant.
We apply the present approach to a gas of point charges. The descriptor is then compared to the long-distance equivariant (LODE) framework Grisafi and Ceriotti (2019) and the MPNN MACE Batatia et al. (2022) for liquid sodium chloride and zirconia.
@@ II.1 Short-Range Descriptors
from Ref. Jinnouchi, Karsai, and Kresse, 2019 that are given there in Eq. (18) and (19) are used for building the short-range descriptors. Here, the vector $\textbf{r}_{jk}$ is the vector pointing from atom $j$ to $k$, $\textbf{r}_{k}-\textbf{r}_{j}$. The set of radial $\chi_{nl}$ and angular $Y_{lm}$ basis functions are used to expand the atomic density in order to obtain rotational and translational invariant expansion coefficients Bartók et al. (2010); Bartók, Kondor, and Csányi (2013). The indices $l$, $m$, and $n$ are the angular, momentum, and radial index, respectively. The parameter $\sigma$ is used to broaden the density distribution. $\iota_{l}$ are the modified spherical Bessel functions of the first kind, and $N_{J}$ is the number of atoms of type $J$. The cutoff function $f_{\text{cut}}$ is used to ensure a smooth decay of $h_{nl}$ to zero at the cutoff radius.
@@ II.2 Long-Range Descriptors
The objective is to calculate a similar set of descriptors where the expansion coefficients however represent the atomic density over long ranges and multiple supercells. The same ansatz as in real space is used, where the expansion coefficients are designed to represent the atomic density around each atom. As before, the descriptors are obtained by projecting the density onto a product basis of radial and angular functions. However, to avoid introducing a cutoff, the corresponding products are evaluated in reciprocal instead of real space:
@@
where G are momentum vectors consistent with the periodic boundary conditions, and $f_{nlm}(\textbf{G})$ are the basis functions in reciprocal space and the other terms represent the broadened atomic density distribution of atom $k$ of type $J$ with positions $\textbf{r}_{k}$ around atom $j$ with positions $\textbf{r}_{j}$. The central atom $j$ may be of any atom type present in the material. The translational invariance is obvious from Eq. (9), as seen from the difference vector $\textbf{r}_{k}-\textbf{r}_{j}$, where $\textbf{r}_{j}$ is the position of the central atom.
@@
The long-range expansion coefficients given in Eq. (9) have the same form as the short-range expansion coefficients given in Eq. (1). Like conventional two- (and three-) body descriptors, they are also atom-centered and invariant under translations and rotations. So for all practical purposes, they behave and can be used like conventional short-range descriptors. For example, the learned regression coefficients are transferable to different unit cells and do not depend on the choice of the Bravais lattice. The only difference is that the evaluation of the expansion coefficients $c_{nlm}^{jJ}$ is performed in reciprocal space. This allows to cover arbitrary interaction ranges, and thus, to encode and measure long-range density fluctuations.
With the appropriate choice of basis functions (see next section), the long-range expansion coefficients can encode the long-range density correlations. This is due to the calculation in reciprocal space, where not only the position of an atom is considered but also all periodic images of this atom. It is important to note that these long-range descriptors do not require a cutoff, but of course, it is possible to impose a cutoff by using the FTs of finite-ranged real-space functions in Eq. (11).
The individual long-range expansion coefficients are then combined into a vector ("long-range descriptor"). In the present work, we mostly use two-body expansion coefficients as shown in Eq. (4) and the corresponding vector $\textbf{X}_{j}^{(2)}$. The three-body long-range coefficients are excluded since the objective here is to describe long-range pairwise interactions. Furthermore, we found no improvements when including long-range three-body terms in any of the systems considered here. This might have many reasons, but we believe that three-body interactions are generally shorter-ranged. Consider, for instance, the case of Van der Waals (vdW) interactions: two-body vdW interactions fall off like $1/r^{6}$, and the three-body terms (Axilrod–Teller) fall off like $1/r^{9}$.
@@ II.3 Long-Range Radial Basis Functions
Spherical harmonics are also used for the angular part of the long-range descriptor. However, spherical Bessel functions cannot be employed to model the infinitely ranged interactions in the radial part, as to maintain a fixed spatial resolution, the number of Bessel functions would need to increase linearly with $R_{\text{cut}}$.
Given the objective of representing electrostatic interactions that decay slowly like $1/r$, where $r$ is the distance between two atoms, it is evident that exponentially decaying functions are a sensible choice. This is because
@@
holds. The first line of Eq. (12) is the Laplace transform of $1$ resulting in $1/r$, while the second line approximates this integral using a quadrature rule. This demonstrates that the integral over all exponentially decaying functions can represent $1/r$, a fact that has been amply used in quantum chemistry to deal with integrals in many-body perturbation theory Häser and Almlöf (1992); Izmaylov and Scuseria (2008); Doser, Lambrecht, and Ochsenfeld (2008); Ayala and Scuseria (1999); Kaltak, Klimes, and Kresse (2014). It is important to understand that the $1/r$-function can be represented very accurately by the sum of exponentials Hackbusch (2019). If both, the decay constants and the weights are optimized independently 6, 8 and 10 functions are sufficient to represent $1/r$ with a maximum absolute error of $10^{-5}$, $10^{-6}$ and $10^{-7}$, respectively, on an interval between $r=[r_{\text{min}},100\cdot r_{\text{min}}]$. More functions are required as the required range increases, but the number of functions increases slowly. It is also found that the optimal decay constants are typically exponentially spaced, with the typical factor between successive decay constants being close to two. We, therefore, expect that if the shortest nearest neighbor distance is around, say, 2 Å, then about 6-8 functions should be sufficient to describe long-range interactions up to typically 200 Å with very high accuracy.
For simplicity, we here define $\zeta_{n}$ on a logarithmic mesh where
@@
and $\zeta_{\text{max}}$ and the scaling constant $s$ are hyperparameters determined numerically. As already noted, this choice was inspired by our experience with many-body applications where a roughly exponential scaling was also found to be optimal Kaltak, Klimes, and Kresse (2014).
@@ II.4 The LODE Implementation
Because of the relation shown in Eq. (14), it is relatively easy, with the density-based model presented above, to implement the LODE descriptors. The only required step is to multiply the density-based expansion coefficients by the negative Coulomb kernel before summing over the reciprocal grid points G. Spherical Bessel functions are used as radial basis functions for this LODE implementation.
As in the original work of Grisafi and Ceriotti, for real materials, the LODE descriptors are combined with short-range descriptors as well, using the approach shown in the following section.
@@ II.5 Combination of Short- and Long-Range Descriptors
To apply the models to realistic materials in which both short-range and long-range interactions are present, the final descriptor vectors X must contain information about both types of interaction. By simply appending the short-range descriptors $\textbf{X}_{\text{sr}}$ and long-range descriptors $\textbf{X}_{lr}$ the final descriptor
@@
As the long-range descriptors are very similar in construction to the short-range descriptors, additional scaling was found to be irrelevant for them. However, to obtain an accurate MLFF involving LODE, we found it crucial to optimize the weight of the LODE descriptors using a hyper-parameter search, as the LODE descriptors have a potentially very different scale as other descriptors.
@@ II.6 The Fitting Process
The fitting process of our model employs the methods of polipy4vasp, the code used in Ref. Schmiedmayer and Kresse, 2024. These methods are implemented analogously to those of the Vienna Ab initio Simulation Package (VASP) Kresse and Furthmüller (1996a, b); Kresse and Joubert (1999) which are presented in Refs. Jinnouchi, Karsai, and Kresse, 2019; Jinnouchi et al., 2019, 2020a.
@@
of local energy contributions $U_{j}$ that depend on the "local" environment around each atom $j$. Kernel regression is used to describe the non-linear dependence of the energy on the descriptors, and a sparse representation is used for the kernel. Either a Gaussian (radial basis function) kernel
@@
In Eq. (18) $\hat{\textbf{X}}$ indicates that the descriptor X is normalized. The index $b$ in both Eq. (17) and Eq. (18) refers to the kernel basis function to which the descriptor of the current central atom $\textbf{X}_{j}$ is "compared".
@@
During training the linear system of equations is calculated via the pseudoinverse of $\bm{\Phi}$ using a singular value decomposition, and singular values smaller than a threshold of $10^{-9}$ times the largest singular value are disposed of.
@@ III Results
For LODE the model with only pairwise descriptors and a body order term of $\nu=1$ is used. This facilitates a comparison with the purely radial description of the density-based long-range descriptors. Furthermore, it was found that the use of three-body terms in our case leads to larger errors compared to the use of only two-body terms. This is likely related to the already discussed two-body nature of long-range Coulomb interactions.